# Advancing Textual Prompt Learning with Anchored Attributes

### v3

@@ 1 Introduction
Vision-Language Models (VLMs) [33, 45, 15, 40, 44, 27, 26], such as CLIP [40] and ALIGN [15], have demonstrated exceptional performance in recent years. These models are trained with a contrastive loss to establish alignment between image and text (category) space. Inspired by the success of NLP [24, 25], prompt learning [16, 61, 60] has emerged as a parameter-efficient tool to adapt powerful VLMs to downstream tasks. Models with a few learnable soft prompt tokens can achieve performance parity with, or even outperform, fully fine-tuned ones [16]. Depending on how the soft prompt tokens are applied, existing methods can be broadly classified into textual-based [61, 60, 18, 19, 29, 51] and visual-based approaches [1, 2, 16, 22, 57]. Among these, the textual-based method is the most fundamental and straightforward, comprising the majority.
@@ 2 Related Work
Prompt Learning for VLMs. Inspired by recent advancements in NLP [24, 25], prompt learning [61, 60, 62, 18, 42, 19, 29, 58] has garnered significant interest among vision researchers aiming to apply these techniques to VLMs [40, 15, 53], such as CLIP. CoOp [61] is the pioneering text-based approach that introduced the concept of using a combination of soft textual tokens and a hard class token as input. Subsequent studies [60, 55, 18, 23, 19, 29] have predominantly followed this textual prompt format. However, this form constrains the soft prompts to align with images within a one-dimensional, predefined category space, limiting their applicability to unknown categories. Therefore, training based on the current text form will be more likely to overfit to known categories, diminishing their zero-shot generalization capability for unknown categories. To address this limitation, several methods have been proposed [55, 17, 62, 19, 29]. For instance, KgCoOp [55] uses hand-crafted hard prompts to regularize learnable soft prompts during training. PromptSRC [19] utilizes CLIP’s [40] original features to regularize the learning of soft prompts for image and text branches. PromptKD [29] utilizes a pre-trained strong teacher model to guide the learning of a student model [13, 28, 54, 30] with learnable prompts. Despite these advancements, none of the above methods address the inherent limitations of the format itself. In this work, we introduce the attribute-anchored textual prompt format for VLMs, which proposes to utilize attributes as a bridge to build more accurate associations between images and unknown categories.
@@ 3.2 Learning soft prompts with universal attributes
Other than placing the hard class token at the end of the text prompt, we can also place it in the middle or front. In Tab. 5, we verify the performance of each position and select the best-performing end position as our default form.
@@ 3.3 Attribute search
Cost Analysis. Unlike traditional Neural Architecture Search (NAS) methods [31, 8, 46], which search computationally expensive network-level parameters, our approach focuses on a lightweight token-level search space. This design makes our method significantly more efficient than previous approaches. In practice, our search converges in approximately 5 epochs, oftern requiring less than 5 minutes on a single A800 GPU. This highlights the practical advantages of our method over traditional parameter architecture search. In addition, search efficiency can be improved by curating a smaller set of base attributes, which narrows the search space.
@@ 4.1 Settings
Attribute Search. We select five independent attributes as the basis in the attribute pool. This results in 31 candidate attribute combinations for the search process. We use ChatGPT-4o for attribute queries. In Tab. 3 reports a subset of the queried attribute bases and the final combination selected by our search. The Appendix provides further details on our search process.
Implementation Details. We evaluate the model performance on 15 popular recognition datasets. We report base and novel class accuracy and their harmonic mean (HM) averaged over 3 runs. The details of each dataset are attached in the Appendix.
@@ 4.5 Further Analysis
By default, the experiments are conducted on the ImageNet. To minimize the influence of other components, we mainly adopt CoOp as the baseline method. Two attributes (color and shape) are used in our ATPrompt. For more experimental results, please refer to the Appendix.
@@
Comparison to Other Attributes. In Tab. 8, we explore the effectiveness of attributes derived through alternative methods, specifically by manually selecting class-irrelevant and common attributes. The results indicate that manually selected irrelevant attributes exhibit comparable performance during training; however, they perform poorly when applied to new categories. This suggests that incorrect attribute tokens cause the soft tokens to develop biased representations, thereby diminishing their zero-shot generalization ability.
@@ 5 Conclusion
Acknowledgments. This work was supported by the National Natural Science Foundation of China under Grant Nos. 62361166670 and U24A20330. This research was also supported by the Young Scientists Fund of the National Natural Science Foundation of China (Grant No.62206134), the Fundamental Research Funds for the Central Universities 070-63253222 and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP). This work was also supported by Shenzhen Science and Technology Program (JCYJ20240813114237048). Computation is supported by the Supercomputing Center of Nankai University (NKSC).
@@
6 Implementation Details
6.1 Dataset
@@ 6.2 Attribute Search
Tab. 11 presents the five attribute bases generated by the LLM, alongside the optimal attribute combination identified after the search. Furthermore, Tab. 13 displays the final weights of all candidate combinations from the search stage on the Caltech-101 dataset.
6.3 Base-to-Novel Generalization
@@
Figure 6: Architectural comparison between CoOp and CoOp+ATPrompt.
CoOp+ATPrompt: Following the baseline, we use a batch size of 32 and an initial learning rate of 0.002. The original paper reports a learnable prompt length of $M=16$ for ResNet-50 but does not specify a length for ViT-B/16. In our setup, we set the sofo token length for both the attribute and class tokens to 2. While the baseline model is trained for 200 epochs, we reduce the training to 100 epochs while maintaining the same cosine decay schedule. Figure 6 illustrates the architectural differences between the original CoOp and CoOp+ATPrompt.
@@
CoCoOp’s original design uses a meta-network to generate offsets for all soft prompt tokens. We retain this meta-network but modify its application: the meta tokens now serve as offsets exclusively for the class soft tokens, $[T_{1}],...,[T_{M}]$, as shown in Fig. 7.
Figure 7: Architectural comparison between CoCoOp and CoCoOp+ATPrompt. In CoCoOp+ATPrompt, meta tokens are only added as offsets to class soft tokens.
@@
The primary architectural modification in MaPLe + ATPrompt concerns the projection mechanism. The original MaPLe framework inputs all textual soft tokens into a projection layer to generate corresponding visual tokens, which are then fused into the image encoder. Our approach, however, selectively inputs only the class soft tokens into this projection layer, while the attribute tokens are preserved without modification. This architectural difference is visualized in Fig. 8.
DePT+ATPrompt: We adopt the baseline’s training configuration: a batch size of 32, a learning rate of 0.0035, a balance weight of $\lambda$=0.7, and a duration of 10 epochs. Our primary configuration for DePT+ATPrompt uses a learnable token length of 4. For datasets with lower complexity, namely Caltech-101, OxfordPets, and StanfordCars, we adjust these parameters, setting the soft token length to 2 and the balance weight to 0.6. The architectural differences between the DePT and DePT+ATPrompt models are detailed in Fig. 9.
Figure 8: Architectural comparison between MaPLe and MaPLe+ATPrompt.
Figure 9: Architectural comparison between DePT and DePT+ATPrompt.
7 Additional Experiments
7.1 Ablation Study
Attribute Order. In the main paper, our experiments confirm that the order of attributes does not significantly impact model performance, with results fluctuating within an acceptable range. Here we provide additional experiments in Tab. 9 to support this observation.
@@ 7.1 Ablation Study
Table 9: Comparison of different attribute orders on ImageNet. Changes in attribute order will not significantly affect model performance.
Attribute Position. We also investigated the impact of attribute token positioning within the prompt. Fig. 10 visualizes the positions tested, and Tab. 10 presents the results. Our findings show that the ”interval” configuration, where attributes are placed between class tokens, yields the best performance.
Figure 10: Comparison of attribute tokens at different positions, taking two attributes as an example.
@@
Table 10: Performance results of attribute tokens at different positions in ATPrompt on ImageNet. The interval version achieves best results.
@@
Table 11: Attribute bases and searched results for each dataset.
Initialization. Baseline methods typically initialize soft tokens using the embeddings of the phrase ”a photo of a.” The inclusion of attribute tokens makes this strategy suboptimal for our method. We instead initialize class soft tokens ($[T_{1}],...,[T_{M}]$) by sampling from a Gaussian distribution with a mean of 0 and a standard deviation of 0.02. As shown in Table 12, this random initialization provides a superior starting point for training.
Attribute Base Novel HM
“a photo of a” 76.40 70.07 73.10
Random Normal Init 76.27 70.60 73.33
Table 12: Comparison of different initialization ways on ImageNet. Random normal initialization performs better.
8 Discussion
9 Limitations and Future Works.
Attributes shape, color, material, function, size
@@ 8 Discussion
Table 13: Output results after 20 epochs of attribute searching on the Caltech101 dataset.

### v4

@@ 1 Introduction
Vision-Language Models (VLMs) [34, 47, 16, 42, 35, 46, 28, 27], such as CLIP [42, 4] and ALIGN [16], have demonstrated exceptional performance in recent years. These models are trained with a contrastive loss to establish alignment between image and text (category) space. Inspired by the success of NLP [25, 26], prompt learning [17, 64, 63] has emerged as a parameter-efficient tool to adapt powerful VLMs to downstream tasks. Models with a few learnable soft prompt tokens can achieve performance parity with, or even outperform, fully fine-tuned ones [17]. Depending on how the soft prompt tokens are applied, existing methods can be broadly classified into textual-based [64, 63, 19, 20, 30, 53] and visual-based approaches [1, 2, 17, 23, 60]. Among these, the textual-based method is the most fundamental and straightforward, comprising the majority.
@@ 2 Related Work
Prompt Learning for VLMs. Inspired by recent advancements in NLP [25, 26], prompt learning [64, 63, 65, 19, 44, 20, 30, 61] has garnered significant interest among vision researchers aiming to apply these techniques to VLMs [42, 16, 56], such as CLIP. CoOp [64] is the pioneering text-based approach that introduced the concept of using a combination of soft textual tokens and a hard class token as input. Subsequent studies [63, 58, 19, 24, 20, 30, 54] have predominantly followed this textual prompt format. However, this form constrains the soft prompts to align with images within a one-dimensional, predefined category space, limiting their applicability to unknown categories. Therefore, training based on the current text form will be more likely to overfit to known categories, diminishing their zero-shot generalization capability for unknown categories. To address this limitation, several methods have been proposed [58, 18, 65, 20, 30]. For instance, KgCoOp [58] uses hand-crafted hard prompts to regularize learnable soft prompts during training. PromptSRC [20] utilizes CLIP’s [42] original features to regularize the learning of soft prompts for image and text branches. PromptKD [30] utilizes a pre-trained strong teacher model to guide the learning of a student model [14, 29, 57, 31] with learnable prompts. Despite these advancements, none of the above methods address the inherent limitations of the format itself. In this work, we introduce the attribute-anchored textual prompt format for VLMs, which proposes to utilize attributes as a bridge to build more accurate associations between images and unknown categories.
@@ 3.2 Learning soft prompts with universal attributes
While this example places the class token at the end, we also evaluate configurations where it is positioned at the front or in the middle. As shown in Tab. 5, the end position yields the best performance and is therefore adopted as our default configuration.
@@ 3.3 Attribute search
Cost Analysis. Unlike traditional Neural Architecture Search (NAS) methods [32, 9, 48], which search computationally expensive network-level parameters, our approach focuses on a lightweight token-level search space. This design makes our method significantly more efficient than previous approaches. In practice, our search converges in approximately 5 epochs, oftern requiring less than 5 minutes on a single A800 GPU for some datasets. In addition, search efficiency can be improved by curating a smaller set of base attributes, which narrows the search space.
@@ 4.1 Settings
Attribute Search. We select five independent attributes as the basis in the attribute pool. This results in 31 candidate attribute combinations for the search process. We use ChatGPT-4o for attribute queries. In Tab. 3 reports a subset of the queried attribute bases and the final combinations selected by our search.
Implementation Details. We evaluate the model performance on 15 datasets. We report base and novel class accuracy and their Harmonic Mean (HM) averaged over 3 runs. The details of each dataset are attached in the Appendix.
@@ 4.5 Further Analysis
By default, the experiments are conducted on the ImageNet. To minimize the influence of other components, we mainly adopt CoOp as the baseline method. Two attributes (color and shape) are used in our ATPrompt.
@@
Comparison to Other Attributes. In Tab. 8, we explore the effectiveness of attributes derived through alternative methods, specifically by manually selecting class-irrelevant and common attributes. It shows that manually selected irrelevant attributes exhibit comparable performance during training; however, they perform poorly when applied to new categories. This suggests that incorrect attribute tokens cause the soft tokens to develop biased representations, thereby diminishing their zero-shot generalization ability.
@@ 5 Conclusion
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant Nos. 62361166670 and U24A20330. This research was also supported by the Young Scientists Fund of the National Natural Science Foundation of China (Grant No.62206134), the Fundamental Research Funds for the Central Universities 070-63253222 and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP). Computation is supported by the Supercomputing Center of Nankai University (NKSC).
@@ 6 Acknowledgments
S1 Implementation Details
S1.1 Dataset
@@
Tab. S4 presents the five attribute bases generated by the LLM, alongside the optimal attribute combination identified after the search. Furthermore, Tab. S5 displays the final weights of all candidate combinations from the search stage on the Caltech-101 dataset.
S1.3 Base-to-Novel Generalization
@@
CoOp+ATPrompt: Following the baseline, we use a batch size of 32 and an initial learning rate of 0.002. The original paper reports a learnable prompt length of $M=16$ for ResNet-50 but does not specify a length for ViT-B/16. In our setup, we set the sofo token length for both the attribute and class tokens to 2. While the baseline model is trained for 200 epochs, we reduce the training to 100 epochs while maintaining the same cosine decay schedule. Figure S1 illustrates the architectural differences between the original CoOp and CoOp+ATPrompt.
@@
CoCoOp’s original design uses a meta-network to generate offsets for all soft prompt tokens. We retain this meta-network but modify its application: the meta tokens now serve as offsets exclusively for the class soft tokens, $[T_{1}],...,[T_{M}]$, as shown in Fig. S2.
Figure S1: Architectural comparison between CoOp and CoOp+ATPrompt.
Figure S2: Architectural comparison between CoCoOp and CoCoOp+ATPrompt. In CoCoOp+ATPrompt, meta tokens are only added as offsets to class soft tokens.
@@
The primary architectural modification in MaPLe + ATPrompt concerns the projection mechanism. The original MaPLe framework inputs all textual soft tokens into a projection layer to generate corresponding visual tokens, which are then fused into the image encoder. Our approach, however, selectively inputs only the class soft tokens into this projection layer, while the attribute tokens are preserved without modification. This architectural difference is visualized in Fig. S3.
DePT+ATPrompt: We adopt the baseline’s training configuration: a batch size of 32, a learning rate of 0.0035, a balance weight of $\lambda$=0.7, and a duration of 10 epochs. Our primary configuration for DePT+ATPrompt uses a learnable token length of 4. For datasets with lower complexity, namely Caltech-101, OxfordPets, and StanfordCars, we adjust these parameters, setting the soft token length to 2 and the balance weight to 0.6. The architectural differences between the DePT and DePT+ATPrompt models are detailed in Fig. S4.
Figure S3: Architectural comparison between MaPLe and MaPLe+ATPrompt.
Figure S4: Architectural comparison between DePT and DePT+ATPrompt.
S2 Additional Experiments
S2.1 Ablation Study
Attribute Order. In the main paper, our experiments confirm that the order of attributes does not significantly impact model performance, with results fluctuating within an acceptable range. Here we provide additional experiments in Tab. S1 to support this observation.
@@
Table S1: Comparison of different attribute orders on ImageNet. Changes in attribute order will not significantly affect model performance.
Attribute Position. We also investigated the impact of attribute token positioning within the prompt. Fig. S5 visualizes the positions tested, and Tab. S2 presents the results. Our findings show that the ”interval” configuration, where attributes are placed between class tokens, yields the best performance.
Figure S5: Comparison of attribute tokens at different positions, taking two attributes as an example.
@@
Table S2: Performance results of attribute tokens at different positions in ATPrompt on ImageNet. The interval version achieves best results.
Initialization. Baseline methods typically initialize soft tokens using the embeddings of the phrase ”a photo of a.” The inclusion of attribute tokens makes this strategy suboptimal for our method. We instead initialize class soft tokens ($[T_{1}],...,[T_{M}]$) by sampling from a Gaussian distribution with a mean of 0 and a standard deviation of 0.02. As shown in Table S3, this random initialization provides a superior starting point for training.
Attribute Base Novel HM
“a photo of a” 76.40 70.07 73.10
Random Normal Init 76.27 70.60 73.33
Table S3: Comparison of different initialization ways on ImageNet. Random normal initialization performs better.
S3 Discussion
S4 Limitations and Future Works.
@@
Table S4: Attribute bases and searched results for each dataset.
Attribute Bases shape, color, material, function, size
@@
Table S5: Output results after 40 epochs of attribute searching on the Caltech101 dataset.